The area of a kite is 384 square feet. One diagonal is three times as long as the other diagonal. Find the lengths of each diagonal.
Answer by Boreal(15235)

Area of kite is pq/2, where p and q are diagonals
q=3p
area is 3p^2/2=384 sq ft
3p^2=768 sq ft
p^2=256 sq ft
p=16 feet.
q=48 feet
Those are the diagonals
Half their product is 384 sq ft.